Hello. I got stumped on this question. I have been trying to complete the square but then I got this question and now I don't know which way to go. Can you help me out? Here is the question:
x^2-16=-6x
Would I add 6x to both sides?

Answer by Alan3354(69443)

x^2-16=-6x
Would I add 6x to both sides?
Why not?
x^2 + 6x - 16 = 0
Add 16
x^2 + 6x = 16
(b/2)^2 = 9
Add 9
x^2 + 6x + 9 = 25
(x+3)^2 = 5^2
etc